OpenStudy (thcsbabygirl1224):

The area of a circle is 78.5 square centimeters, and a subtending arc on the circle has an arc length of 6pi. The estimated value of pi is 3.14. The measure of the angle subtended by the arc is °.

OpenStudy (thcsbabygirl1224):

@adilalvi @AloneS. sorry i keep tagging you i just truely understand this stuff

OpenStudy (alones.):

A = pi r^2 78.5 = pi r^2 r = sqrt(78.5/pi) arc length in radians 6 pi cm / (sqrt(78.5/pi) cm/rad) = 3.77 rad or 3.77 rad / (pi rad / 180°) = 120° i assumed that you meant 6 pi cm for arc and the central angle is equal to subtended arc measure

OpenStudy (thcsbabygirl1224):

i did thank you you explain it to where i understand

OpenStudy (alones.):

Yep!